OpenStudy (anonymous):

How can we use transformations to find the vertical asymptote of a logarithmic functions. I can't figure this out.

OpenStudy (anonymous):

@JDETSC

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Be more specific, not really sure what you're asking. Post an example.

OpenStudy (anonymous):

idk that's the question my teacher is asking, that's why I'm stuck it makes no sense.

OpenStudy (agent0smith):

If you mean something like y = log(x+5) then that is the graph of y = logx, but shifted 5 units left.... when means the vertical asymptote shifts 5 units left. Do you know the vertical asymptote of y=logx?

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

I guess that makes sense although that's not how he worded the question that's probably what he's asking. Thank you.

OpenStudy (agent0smith):

Well idk what to make of the actual question. You could make it general by saying that for y=log(x - a), the vertical asymptote is x=a.

OpenStudy (anonymous):

okay, how do I give a medal or something because I registered just for this question.

OpenStudy (anonymous):

nvm